# A Survey of AI Methods for Geometry Preparation and Mesh Generation in Engineering Simulation

### v1

@@ Abstract
Artificial intelligence is beginning to ease long-standing bottlenecks in the CAD-to-mesh pipeline. This survey reviews recent advances where machine learning aids part classification, mesh quality prediction, and defeaturing. We explore methods that improve unstructured and block-structured meshing, support volumetric parameterizations, and accelerate parallel mesh generation. We also examine emerging tools for scripting automation, including reinforcement learning and large language models. Across these efforts, AI acts as an assistive technology, extending the capabilities of traditional geometry and meshing tools. The survey highlights representative methods, practical deployments, and key research challenges that will shape the next generation of data-driven meshing workflows.
@@ 1 Introduction
High-fidelity simulation is now a cornerstone of engineering and scientific workflows. Yet, getting from a 3D model to a simulation-ready mesh remains one of the most time-consuming and frustrating steps in the process. Engineers still spend countless hours identifying parts, simplifying geometry, setting up meshing parameters, and checking mesh quality. These steps are often repeated across similar models. Despite decades of progress in geometry kernels and meshing algorithms, this part of the workflow remains surprisingly manual and error-prone.
At the 2024 International Meshing Roundtable, a panel of researchers and practitioners, including several authors of this paper, discussed whether artificial intelligence could help. The answer was not that AI will replace traditional methods. Rather, AI might help amplify them by learning from past decisions, catching problems early, and automating tedious steps. In this view, AI is a powerful assistant that complements expert knowledge instead of replacing it.
Artificial intelligence (AI) and machine learning (ML) are broad terms that describe systems that learn from data. In recent years, these tools have made significant progress in areas such as image recognition, language understanding, and robotics. They are now beginning to show promise in simulation workflows. Applications include identifying parts in CAD assemblies, predicting where meshes might fail, guiding how geometry should be simplified, and even generating scripts to automate routine tasks.
To help readers navigate this growing space, we present a taxonomy in Table 1. It summarizes the main types of AI and machine learning methods being used in model preparation and meshing. Each category includes a brief description, examples of how it is applied, and representative citations. For readers who are new to this area, the table provides both a quick reference and a roadmap for the rest of the paper.
Table 1: Taxonomy of AI/ML methods applied to CAD model preparation and meshing workflows, showing descriptions, referenced sections, and representative citations.
@@
Classical Machine Learning Rule-based statistical learning (SVMs, decision trees, ensembles, surrogates) 3D Part Classification; Mesh Quality Prediction; Defeaturing; Parallel Mesh Generation [IpRegli2006, Jayanti2006, Breiman2001, OwenIMR2019, OwenOSTI2022, OwenIMR2023, Parry2025, ParryIMR2023, sklearn, gamatie2019, ipek2005, garner2025numa, garner2025dist, tsolakis2024]
@@
Deep Learning – Geometric Models Neural networks specialized for geometry (CNNs, GNNs, Transformers, hybrids) 3D Part Segmentation; Mesh Quality Prediction (baseline); Volumetric Parameterizations; Scripting Automation (graph encodings) [lambourne2021brepnet, dupont2022cadops, cha2023machining, lee2023brepgat, wu2024aagnet, zheng2025sfrgnn, park2025dfgat, zhang2024brepmfr, dai2025brepformer, jayaraman2021uv, wang2023hybrid, uwimana2025segmentation, yang2025hybrid, DeepShapeCode, zhang2024brep2seq, wu2020comprehensive]
@@
Generative & Representation Learning Generative shape representations (implicit SDFs, parametric splines, hybrids, diffusion models) Unstructured Meshing (geometry construction); Volumetric Parameterizations (DL-Polycube, DDPM-Polycube, neural integral mapping) [park2019deepsdf, mescheder2019occupancy, peng2020convolutional, mildenhall2021nerf, gropp2020implicit, larios2021deep, novello2022exploring, dong2024neurcadrecon, dong2025neurcross, sharma2020parsenet, williamson2025neural, maruani2024ponq, he2025sparseflex, liao2018deep, chen2022neural, sundararaman2024self, yu2025dlpolycube, yu2025DDPM, Zhan2025]
@@
Reinforce-ment Learning (RL) Agents learn multi-step geometry/meshing actions via trial-and-error Defeaturing (future RL); Unstructured Meshing (quad, Delaunay); Block Structured Meshing (2D, 3D); Volumetric Parameterizations (sequence optimization); Scripting Automation (block decomposition, thin volume reduction, CAD scripting); Language-Driven Scripting (integration with RL) [pan2023reinforcement, TONG2023102109, Thacher2025, diprete2024reinforcement, haarnoja-2018-sac, Patel2021AutoHex, Zhang2025CADRLHex, owen2024cad, zhang2025reinforcement]
@@
Large Language Models (LLMs) Text-to-code synthesis, multimodal assistants, agentic AI (LLM+RL) Language-Driven Scripting; Scripting Automation (integration with GNN/RL) [zhao2023survey, R1, R4, R13, N1, R5, R6, R14, R18, R26, N2, zhang2024brep2seq, zhang2024brepmfr, wu2020comprehensive, fang2025selfevolving]
The range of applications is broad. Classical machine learning methods support tasks such as part classification and mesh quality prediction. Deep learning models designed for geometric data are being used for segmentation, defeaturing, and volumetric parameterization. Reinforcement learning is being tested for guiding multi-step decisions in meshing workflows. Large language models are beginning to assist with scripting, automation, and natural language interaction with CAD tools.
@@
Many of these efforts are still in early stages, but the momentum is clear. AI is beginning to reshape not only how meshes are generated, but also how geometry is prepared, validated, and integrated into simulation pipelines. This paper surveys the current landscape, highlights where AI is making a difference, and outlines the challenges that remain. These include the need for curated datasets, standardized benchmarks, better representations for CAD geometry, and intelligent systems that can support analysts across the entire modeling workflow.
@@
Traditional non-AI approaches rely on manual tagging or rule-based feature recognition. These methods can work for narrowly defined classes but tend to be brittle across diverse CAD sources, modeling conventions, tolerances, and geometric noise, and they scale poorly as category counts and edge cases grow [Jayanti2006, IpRegli2006]. In contrast, AI can learn invariants and decision boundaries from data, tolerating variability in geometry and topology while scaling to many classes. The broader literature spans point-cloud, volumetric, image, and graph-based representations [BRepNet]. Recent B-rep graph methods strengthen the direct use of CAD topology for recognition and segmentation [Colligan2022, Lee2023].
Formally, part classification is cast as supervised learning. Given a training set $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$, with feature vectors $\mathbf{x}_{i}$ and labels $\mathbf{y}_{i}$, the objective is to learn a function $f$ such that $\hat{\mathbf{y}}=f(\mathbf{x})$. Foundational work includes curvature-based descriptors with SVMs for manufacturing-process discrimination and early shape benchmarks for engineering parts [IpRegli2006, Jayanti2006]. Deep-learning approaches for CAD classification followed [Qin2014], while graph and B-rep networks now operate directly on faces, edges, and coedges [BRepNet, Colligan2022, Lee2023]. In applied workflows, Owen et al. compute features from curves, surfaces, and volumes available via a modeling kernel [OwenIMR2023, ACIS], and train ensembles of decision trees, specifically Random Forests [Breiman2001] implemented in scikit-learn [sklearn]. Labels are curated from large single-part repositories such as GrabCAD [GrabCAD] and ABC [Koch2019], with additional pre-categorized corpora like MCB [MCB2020]. Comparative studies in this context show that tree ensembles provide strong accuracy, fast inference, and interpretable feature importances relative to neural baselines on tabular B-rep feature sets [OwenIMR2023, OwenOSTI2022].
Methods and implementation have matured into deployable tools. The Cubit® Geometry and Meshing Toolkit integrates part classification and reduction features to drive predefined idealization recipes for meshing and boundary condition setup [CubitWebsite, OwenIMR2023, OwenIMR2019]. Altair HyperMesh extends this trend with AI-driven shape recognition (ShapeAI) that automatically identifies and groups recurring parts such as bolts and fasteners, representing one of the first production uses of ML in FEA preprocessing [AltairHyperMesh, AltairShapeAI]. Siemens has likewise incorporated machine learning into Simcenter NX through shape recognition and similarity search, enabling automated grouping and selection of geometrically related parts [SimcenterAIWorkflow, SiemensNXShapeRecognition]. At present, however, only a handful of commercial preprocessors have deployed such AI capabilities, underscoring both their promise and the still-limited extent of adoption across the broader simulation tool ecosystem.
@@
Segmentation of CAD models refers to the task of assigning semantic labels to regions within a part, typically at the face level. While part classification applies a single label to an entire component (e.g., “bolt” or “flange”), segmentation provides finer detail, identifying geometric features such as holes, fillets, bosses, or functional surfaces. This granularity is essential for tasks in simulation, manufacturing, and process planning, where different regions of a part may require distinct treatment in meshing, defeaturing, or boundary condition application.
Traditionally, segmentation has relied on rule-based systems that detect features using geometric heuristics like curvature, convexity, and adjacency. These approaches can succeed in narrow, well-defined domains, but often struggle to generalize across diverse CAD sources and modeling practices. Manual segmentation, while accurate, is labor-intensive and impractical for large-scale workflows.
Recent advances in deep learning have enabled more robust and scalable segmentation by learning directly from the geometry and topology of B-rep models. Neural architectures such as convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers are increasingly adapted to CAD representations. These methods can recognize feature patterns without requiring handcrafted rules, generalize across datasets, and support efficient inference in production workflows.
A number of techniques have emerged to exploit B-rep structure. BRepNet [lambourne2021brepnet] introduced the first convolutional framework operating on directed coedges, enabling accurate face labeling. Extensions such as CADOps-Net [dupont2022cadops] predict both feature types and their construction sequences, while Cha et al. [cha2023machining] apply segmentation to machining operations and group coherent regions via connected components. For deployment in constrained environments, lightweight variants based on pruning [gkrispanis2025towards] and transfer learning [gkrispanis2024enhancing] have been proposed.
Graph-based approaches further generalize segmentation by representing faces and edges as attributed graphs. BRepGAT [lee2023brepgat] applies attention-based message passing, and AAGNet [wu2024aagnet] integrates geometric features into an adjacency graph with multi-task learning. Other variants such as SFRGNN-DA [zheng2025sfrgnn] and DFGAT [park2025dfgat] target domain adaptation and custom descriptors for design feature recognition. Meanwhile, transformer-based methods like BrepMFR [zhang2024brepmfr] and BRepFormer [dai2025brepformer] capture both local and global patterns across B-rep structures, enabling effective transfer from synthetic to real-world data.
Hybrid models combine the strengths of different representations. UV-Net [jayaraman2021uv] operates on UV parameter domains using a CNN-GNN hierarchy. DeepFeature [wang2023hybrid] merges rule-based reasoning with learned embeddings, and MeshCAD-Net [uwimana2025segmentation] fuses BRepNet and MeshCNN features for multi-modal learning. Other systems [yang2025hybrid] use primitive decomposition alongside GNNs to manage complex, interactive feature sets.
While most systems remain in the research or prototype stage, early implementations are emerging. BRepNet has been open-sourced and integrated into academic toolkits, and hybrid pipelines show promising results on synthetic benchmarks and curated real-world datasets. Broader deployment in commercial CAD or meshing software is still limited, but progressing.
Looking forward, segmentation will play a foundational role in automating CAD model preparation. It enables fine-grained control over mesh resolution, defeaturing, and physics-based modeling decisions. Future work includes combining segmentation with simulation-aware labels, improving generalization across CAD systems, and integrating segmentation tightly with downstream AI-driven tools for classification, meshing, and analysis.

### v2

@@ Abstract
Artificial intelligence is beginning to reduce the manual effort in the CAD-to-mesh pipeline. Written for meshing and geometry practitioners with limited AI background, this survey organizes recent work by workflow step. We cover part classification and segmentation, mesh quality prediction, and defeaturing. We review AI guidance for unstructured meshing, block-structured meshing in 2D and 3D, and volumetric parameterization, including reconstruction from implicit or sampled geometry. We also discuss parallel mesh generation and scripting automation via reinforcement learning and large language models. Across these topics, AI complements established geometry and meshing algorithms rather than replacing them. We conclude with practical lessons and open challenges in data, benchmarks, and trustworthy integration.
@@ 1 Introduction
High-fidelity simulation is now a cornerstone of engineering and scientific workflows. Yet, getting from a 3D model to a simulation-ready mesh remains one of the most time-consuming and frustrating steps in the process. Engineers still spend countless hours identifying parts, simplifying geometry, setting meshing parameters, and checking mesh quality. Many of these decisions are repeated across similar models, and small differences in geometry or intent can trigger large changes in meshing behavior. Even with mature geometry kernels and meshing algorithms, the day-to-day workflow can feel more like careful craftsmanship than a reliable pipeline.
This is not because the underlying algorithms are weak. It is because meshing sits at the intersection of geometry, analysis intent, solver requirements, and practical constraints. Much of what makes an expert effective is not a single technique, but the ability to choose among many options, to anticipate failure modes, and to apply a consistent strategy across a family of parts. That expertise is valuable, but it is often captured informally: in tribal knowledge, in scattered scripts, and in decisions that are hard to reproduce when the team changes or the model evolves.
At the 2025 International Meshing Roundtable, a panel of researchers and practitioners, including several authors of this paper, discussed whether artificial intelligence could help. The most useful framing was not replacement. The opportunity is assistance. AI can learn from prior examples, flag likely issues earlier, and automate repetitive steps, while conventional geometry processing and meshing remain responsible for robustness and guarantees. In other words, AI can help scale expert practice without pretending that meshing is a solved problem that can be handled by a single black box.
This paper is written for readers who are comfortable with geometry and meshing, but who may have little or no background in modern AI or machine learning. The goal is to make the landscape approachable, to connect each AI idea to a familiar meshing pain point, and to be clear about where the technology is already useful and where it still falls short. Throughout, AI and machine learning refer to data-driven methods that extract patterns from previous models and workflows, then use those patterns to make predictions or recommend actions. The details of specific algorithms matter, but the key question is simpler: what part of the meshing workflow becomes faster, more consistent, or easier to automate when experience is captured in data?
To help navigate the space, Table 1 provides a taxonomy of the main families of AI methods that appear in geometry processing and mesh generation, along with where they show up in this paper. The table is meant as a roadmap rather than a prerequisite. Readers can treat it as a guide to the kinds of problems each method is being used to solve.
Table 1: Taxonomy of AI and machine learning methods used in geometry processing and mesh generation workflows.
@@
Classical Machine Learning Rule-based statistical learning (SVMs, decision trees, ensembles, surrogates) [Hastie2009ESL, Cortes1995SVM, Breiman2001] 3D Part Classification; Mesh Quality Prediction; Defeaturing; Parallel Mesh Generation [IpRegli2006, Jayanti2006, sklearn, OwenIMR2019, OwenOSTI2022, OwenIMR2023, ParryIMR2023, Parry2025, gamatie2019, ipek2005, garner2025numa, garner2025dist]
@@
Deep Learning - Geometric Models Neural networks specialized for geometry (CNNs, GNNs, Transformers, hybrids) [LeCun2015DeepLearning, Bronstein2017GeometricDL, Vaswani2017Attention] 3D Part Segmentation; Mesh Quality Prediction (GNN baselines); Volumetric Parameterizations; Scripting Automation (graph encodings) [lambourne2021brepnet, dupont2022cadops, lee2023brepgat, wu2024aagnet, zheng2025sfrgnn, zhang2024brepmfr, dai2025brepformer, uwimana2025segmentation, jones2023self, hanocka2019meshcnn, Pfaff2021MeshGraphNets, SpraveDrescher2021, Wang2022MeshQualityGNN, jayaraman2021uv, wang2023hybrid, zhang2024brep2seq, wu2020comprehensive]
@@
Generative & Representation Learning Generative shape representations (implicit SDFs, parametric splines, hybrids, diffusion models) [Kingma2014VAE, Goodfellow2014GAN, Ho2020DDPM] Unstructured Meshing (geometry construction); Volumetric Parameterizations (DL-Polycube, DDPM-Polycube, neural integral mapping) [park2019deepsdf, mescheder2019occupancy, gropp2020implicit, peng2020convolutional, mildenhall2021nerf, larios2021deep, novello2022exploring, dong2024neurcadrecon, dong2025neurcross, zhang2023neurogf, liu2020neural, sharma2020parsenet, williamson2025neural, maruani2024ponq, he2025sparseflex, liao2018deep, chen2022neural, sundararaman2024self, DeepShapeCode, yu2025dlpolycube, yu2025DDPM, Zhan2025]
@@
Reinforce-ment Learning (RL) Agents learn multi-step geometry and meshing actions via trial-and-error [SuttonBarto2018RL, Mnih2015DQN, Haarnoja2018SAC] Defeaturing (future RL); Unstructured Meshing (quad, Delaunay); Block Structured Meshing (2D, 3D); Volumetric Parameterizations (sequence optimization); Parallel Mesh Generation (AMR control); Scripting Automation (block decomposition, thin volume reduction, CAD scripting); Language-Driven Scripting (integration with RL) [pan2023reinforcement, tong2023srl, Thacher2025, diprete2024reinforcement, haarnoja-2018-sac, Patel2021AutoHex, Zhang2025CADRLHex, Gatti2022RLGNN, Yang2023RLAMR, owen2024cad, zhang2025reinforcement]
@@
Large Language Models (LLMs) Text-to-code synthesis, multimodal assistants, and agentic workflows (LLM+tools, LLM+RL) [Vaswani2017Attention, Brown2020GPT3, Ouyang2022InstructGPT] Language-Driven Scripting; Scripting Automation (integration with geometric encoders and RL) [zhao2023survey, R1, R2, R13, R14, R18, R26, N2, R9, R10, R12, R23, R25, fang2025selfevolving, Menon2026MultiAgentISC]
The range of applications is broad, but the motivation is consistent. Some methods help a workflow “understand” geometry in ways that support downstream decisions, such as recognizing parts, identifying functional regions, or separating features from context. Other methods help predict outcomes, such as whether a set of meshing choices is likely to produce poor elements or whether a local region is likely to cause failure. Still others are aimed at automation, where the goal is to turn an expert sequence of steps into something that can be repeated reliably, either as a learned policy or as generated scripts that drive existing CAD and meshing tools.
The remainder of the paper follows the workflow from geometric organization to mesh generation and automation. We first discuss part classification and segmentation, because these tasks often determine what information is available to drive meshing choices. We then survey mesh quality prediction and defeaturing, where AI can reduce costly iterations by warning about problems and suggesting targeted changes. Next, we cover unstructured and block-structured meshing, including recent work that uses learning to guide meshing decisions or to construct intermediate representations that enable structured discretizations. We also discuss volumetric parameterization methods, which sit at a key junction between geometric mappings and structured mesh generation. Finally, we address parallel mesh generation and automated scripting, including reinforcement learning and large language models that aim to make workflow automation more accessible to analysts.
@@
Many of these efforts are still in early stages, but the momentum is clear. AI is beginning to reshape not only how meshes are generated, but also how geometry is prepared, validated, and integrated into simulation pipelines. At the same time, practical adoption highlights challenges that appear repeatedly across the literature: the need for curated datasets and shareable benchmarks, reliable representations of CAD geometry that preserve intent, and evaluation criteria that connect algorithmic success to simulation outcomes. This review highlights where AI is already making a difference, and it outlines what must improve for these tools to become dependable partners in everyday meshing practice.
@@
Early work approached part classification with manual tagging and rule-based feature recognition. These systems encode expert heuristics about geometry or manufacturing features and can work for narrow classes, but they are brittle across diverse CAD sources, modeling conventions, and tolerances, and they scale poorly as category counts grow [IpRegli2006, Jayanti2006].
The authors of [MCB2020] organize whole-part classification by data representation. In the multi-view image family, a CAD model is rendered from several viewpoints and the view set is classified; representative methods include MVCNN and RotationNet [Su2015MVCNN, Kanezaki2018RotationNet]. In the volumetric family, tessellated solids are voxelized and processed with 3D CNNs; common baselines include VoxNet and VRN [Maturana2015VoxNet, Brock2016VRN]. In the point family, methods learn directly from sampled points and often excel when local geometric detail matters; examples include PointNet++, PointCNN, and SpiderCNN [Qi2017PointNetPlusPlus, Li2018PointCNN, Xu2018SpiderCNN]. These representation choices benefit from mature vision toolchains and large public datasets and they perform well for broad shape recognition, yet they may discard exact boundary information and can be sensitive to viewpoint, grid resolution, or sampling density [MCB2020]. For simulation-driven workflows, these limitations motivate methods that keep the CAD boundary model intact.
CAD-native approaches preserve the boundary representation and map cleanly to idealization, meshing, and boundary-condition setup. In addition to supervised labeling, B-rep structure can also be used to build relational understanding at the assembly scale without committing to a fixed label set, which complements intent-centric CAD-CAE integration efforts based on explicit knowledge representations [boussuge2019ontology]. Boussuge et al. use tensor factorisation over geometric and topological relations to support similarity-based retrieval and hierarchical grouping of entities and components in CAD assemblies [boussuge2022tensorfactorisation]. While this is not semantic part classification, it can reduce analyst effort by organizing large assemblies into consistent groups that can later be mapped to task-specific categories when needed.
A practical CAD-native path computes tabular attributes from curves, surfaces, and volumes provided by a modeling kernel [ACIS] and then trains a classifier such as a Random Forest [Breiman2001] using scikit-learn [sklearn]. Typical signals include counts by analytic face type and summary ratios over area, curvature, and edge or vertex statistics [OwenIMR2023]. This design aligns with the cues analysts use, yields readable importance scores [Breiman2001], supports incremental updates with user-labeled examples [OwenIMR2023], and delivers strong accuracy with thousands of parts at assembly-scale inference speed [OwenIMR2023, OwenOSTI2022].
Formally, part classification can be cast as a supervised learning problem. Given a training set $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$, with feature vectors $\mathbf{x}_{i}$ that characterize each part and labels $\mathbf{y}_{i}$ that define the category, the objective is to learn a function $f$ such that $\hat{\mathbf{y}}=f(\mathbf{x})$. Early work used handcrafted descriptors with SVMs and established engineering benchmarks [IpRegli2006, Jayanti2006]. Deep learning approaches for CAD classification followed [Qin2014]. Recent studies also explore whole-part classification from B-rep-derived graphs that retain exact boundary semantics [Roj2022CADGraphClassification, Mandelli2022CADGNN]. Labels are commonly drawn from single-part repositories and curated corpora such as GrabCAD, ABC, and MCB [GrabCAD, Koch2019, MCB2020]. Comparative results show that tree ensembles remain competitive baselines on tabular B-rep features, offering fast and interpretable predictions suitable for deployment [OwenIMR2023, OwenOSTI2022].
Methods and implementation have matured into deployable tools. The Cubit Geometry and Meshing Toolkit integrates part classification and reduction features to drive predefined idealization recipes for meshing and boundary condition setup for fasteners [CubitUserDoc1706_2025, OwenIMR2023, OwenIMR2019]. In a typical workflow, a complex assembly is first classified, the user reviews the resulting category assignments in the GUI, and then selects one or more categories (for example, bolts) to apply an appropriate recipe. These recipes encapsulate common actions for simulation preparation, including geometric transforms, meshing, decomposition, and boundary condition application. In Cubit, users can also retrain in situ by selecting new parts for categorization, adding new categories when needed, and persisting the resulting training data for future sessions; a small set of common categories is shipped by default and can be expanded as users contribute examples [OwenIMR2023].
Building on this pattern, other commercial preprocessors are adding related capabilities. Altair HyperMesh offers AI-driven shape recognition (ShapeAI) that identifies and groups recurring parts such as fasteners [AltairHyperMesh, AltairShapeAI], and Siemens Simcenter NX incorporates shape recognition and similarity search to enable automated grouping and selection of geometrically related parts [SimcenterAIWorkflow, SiemensNXShapeRecognition]. Adoption remains early, which underscores both the promise of these approaches and the still-limited extent of deployment across the broader simulation tool ecosystem.
@@
Segmentation of CAD models assigns semantic labels to regions of a part, most often at the face level. In contrast to part classification, which assigns a single label to a whole component (for example, “bolt” or “flange”), segmentation identifies functional and geometric regions such as holes, fillets, chamfers, ribs, bosses, and contact or load-bearing surfaces. This finer structure is central to simulation workflows because it enables region-aware meshing, selective defeaturing, and reliable application of boundary conditions and material models. In practice, segmentation is also closely related to CAD feature recognition, especially when labels correspond to modeling operations or machining features.
Classical segmentation systems rely on rule-based feature detection using curvature, convexity, adjacency, and other geometric heuristics. These approaches can be effective within a narrow modeling style, but they tend to be brittle when applied across different CAD kernels, tolerance conventions, and feature interactions. Manual segmentation remains accurate, but it does not scale to the volume and iteration rate of modern engineering pipelines. Learning-based methods address these limitations by training models to recognize repeatable geometric and topological patterns directly from B-rep structure, and by producing labels with consistent semantics that can drive downstream automation.
Segmentation benchmarks differ not only in size, but also in what the labels represent. A useful split is between design-oriented datasets, where labels reflect modeling operations and sometimes construction steps, and manufacturing-oriented datasets, where labels follow machining features and other process-relevant attributes.
Design-oriented datasets align naturally with design intent. The Fusion 360 Gallery Segmentation dataset established a standard evaluation setting for face-level labels tied to common construction operations [lambourne2021brepnet]. These operation-driven labels provide structured supervision that is meaningful for downstream automation. CC3D-Ops extends this style by adding construction step annotations that group faces created by the same operation, which supports sequence-aware reasoning [dupont2022cadops].
Manufacturing-oriented datasets align with process planning, and they also translate naturally to simulation preparation because machining features often define regions that require controlled mesh sizing, contact treatment, or localized refinement. MFCAD and MFCAD++ provide face-level machining feature labels, with MFCAD++ increasing geometric complexity and including more challenging feature interactions [cao2020graph, colligan2022hierarchical]. More recent datasets add richer supervision, including instance grouping and bottom-surface identification, which makes the output closer to the structured selections engineers perform in practice. MFInstSeg is a representative example that supports semantic segmentation, instance segmentation, and bottom-surface labeling [wu2024aagnet].
Most modern CAD segmentation pipelines can be grouped by how they represent B-rep structure and how they propagate information across it. The common thread is to combine local geometric cues, such as surface type and convexity, with broader context implied by topology and feature interactions.
A widely cited baseline for direct B-rep learning is BRepNet [lambourne2021brepnet]. Its key idea is to use directed coedges to define neighborhoods for topology-aware convolutions, allowing the network to update features on faces, edges, and coedges while respecting the B-rep structure. This topology-native design has been influential because it provides a clean way to combine geometry and connectivity without flattening the model into an unstructured point representation. Several extensions build on this foundation. CADOps-Net [dupont2022cadops] jointly predicts operation types and construction steps, which helps connect segmentation outputs to design history. In machining-oriented settings, face labels can be grouped into coherent feature instances using connected component analysis after per-face classification [cha2023machining]. For deployment constraints, pruning and transfer learning have been used to reduce computational cost and improve adaptation to new data distributions while retaining much of the baseline performance [gkrispanis2025towards, gkrispanis2024enhancing].
Graph-based approaches represent a B-rep as an attributed adjacency graph whose nodes correspond to faces (and sometimes edges), and whose edges encode topological relations. This representation makes it natural to use graph neural networks (GNNs) that propagate information across neighbors while learning which relationships matter most. BRepGAT [lee2023brepgat] uses attention-based message passing to weight influential neighbors, which can help when local evidence is ambiguous. AAGNet [wu2024aagnet] extends this direction with multi-task learning, predicting semantic labels alongside instance groupings and bottom-surface identification. These multi-task outputs align closely with what engineers need for downstream selection and automation, because they recover both feature type and feature extent in a form that can be consumed by rule-based or learned workflows. Domain shift remains a practical obstacle, since models trained on curated benchmarks can degrade on industrial parts. Approaches such as SFRGNN-DA [zheng2025sfrgnn] target cross-domain robustness, while self-supervised and few-shot strategies aim to reduce dependence on dense manual labels when adapting to new feature libraries or conventions [jones2023self].
Transformer-based models have also been introduced to capture broader context across a part, which is helpful for interactive features, repeated patterns, and cases where local geometry is not sufficient to determine the correct label. BrepMFR [zhang2024brepmfr] combines transformer-style attention with graph structure and uses domain adaptation to improve transfer from synthetic to real CAD models. BRepFormer [dai2025brepformer] similarly propagates information through transformer layers while incorporating topological cues, and it has demonstrated strong performance on multi-task benchmarks such as MFInstSeg [wu2024aagnet]. In parallel, hybrid and multi-modal models combine complementary representations to balance robustness and expressiveness, for example by pairing UV-domain surface encodings with graph propagation [jayaraman2021uv], mixing learned inference with rule-based decomposition for highly interactive features [wang2023hybrid], or fusing B-rep and mesh features [uwimana2025segmentation, hanocka2019meshcnn]. Other directions address practical failure modes such as small-face underrepresentation through sampling-aware strategies and loss designs that improve fine-detail reliability [zhang2023extending].
Most CAD segmentation systems remain research-stage, and broad commercial deployment is still limited. Even so, open-source baselines and better benchmarks are steadily improving reproducibility and lowering the barrier to integration. For simulation and meshing, the most impactful next steps are likely to be labels that are directly aligned with analysis decisions, such as identifying regions that require boundary-layer resolution, separating contact from non-contact surfaces, and predicting defeaturing candidates conditioned on target physics. Equally important is robust generalization across CAD kernels and modeling practices, since engineering pipelines depend on consistent behavior rather than best-case accuracy on a single benchmark. As these gaps close, segmentation can become a foundation for more automated CAD preparation, linking geometry understanding to defeaturing, mesh generation, and analysis setup.
Contributed by N. Lei, X. Gu